I do not understand how to work this problem: y=x^3-8x. I know that the y-intercept is (0,0), but I can not figure out how to find the x-intercepts.
Answer by stanbon(75887)

y=x^3-8x. I know that the y-intercept is (0,0), but I can not figure out how to find the x-intercepts.
----------------
To find the x-intercepts, let y=0.
x^3-8x=0
x(x^2-8)=0
x=0 or x^2-8=0
x=0 or x=2sqrt2 or x=-2sqrt2
==================================
Cheers,
Stan H.